OpenStudy (anonymous):

Indicate the equation of the line that is the perpendicular bisector of the segment with endpoints (4, 1) and (2, -5).

OpenStudy (anonymous):

HELP ASP

OpenStudy (anonymous):

first you create equation of line with given points.

OpenStudy (anonymous):

I just need the answer right now

OpenStudy (anonymous):

then line perpendicular to it have a slope in the form . m*m'=-1. so you can proceed further

OpenStudy (eric_d):

y=-3x+c @Joseph91

OpenStudy (anonymous):

i didn't tried the question.But i have given some ideas which will help you to solve the question.

OpenStudy (eric_d):

@mathmale

OpenStudy (anonymous):

It has + and the = as the preset in the beginning

OpenStudy (anonymous):

Here is the question

OpenStudy (anonymous):

@eric_d

OpenStudy (eric_d):

hang on

OpenStudy (eric_d):

first find midpoint

OpenStudy (eric_d):

then find the gradient, m

OpenStudy (anonymous):

eric i will give you another medal plus a testimonial

OpenStudy (eric_d):

midpoint= (3,-2)

OpenStudy (eric_d):

gradient,m= -5-1/2-4=3

OpenStudy (anonymous):

Im in a test thats why i need the answer

OpenStudy (eric_d):

since it's perpendicular bisector, m=1/3

OpenStudy (eric_d):

so,

OpenStudy (eric_d):

y=m(x-a)+b

OpenStudy (eric_d):

y=1/3(x+3)-2

OpenStudy (eric_d):

y=1/3x-1

OpenStudy (eric_d):

that's the final answer @fail

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):

Its not right but its ok

OpenStudy (eric_d):

but that's what I got

OpenStudy (anonymous):

@eric_d did you see the photo that's how its suppose to be

OpenStudy (eric_d):

yup I'm nt very sure with that Try tagging others but that's what I got

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@sweetburger